∫lnx dx

∫lnx dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int lnxdx
I = \int x^{0}lnxdx
Apply ILATE Rule
I = xlnx - \int dx
I = xlnx - x + c
I = x(lnx - 1) + c